# Evidential Trust-Aware Model Personalization in Decentralized Federated Learning for Wearable IoT

### v1

@@ VII-A1 Datasets
  • UCI HAR [anguita2013public]: Smartphone accelerometer and gyroscope data from 30 subjects performing 6 activities (walking, walking upstairs, walking downstairs, sitting, standing, lying). Contains 10,299 samples with 561 extracted features.
  • PAMAP2 [reiss2012introducing]: IMU data from 9 subjects performing 12 activities including household and exercise tasks. We extract 40 features per sliding window from three body-worn sensors.
  • PPG-DaLiA [reiss2019deep]: Photoplethysmography and accelerometer signals from 15 subjects under real-life conditions. We formulate activity classification from 8 activity types with time-frequency features.
@@ VII-A2 Baselines
  • FedAvg [mcmahan2017communication]: Standard federated averaging without Byzantine resilience or personalization mechanisms.
  • BALANCE [fang2024byzantine]: Distance-based filtering with adaptive thresholds for Byzantine resilience.
  • Sketchguard [rangwala2025sketchguard]: Count-Sketch compression with cosine similarity filtering for communication-efficient Byzantine resilience.
  • UBAR [guo2021byzantine]: Two-stage filtering combining distance-based screening with loss-based evaluation.
@@ VII-A3 Implementation Details
We deploy 30 nodes in a fully-connected topology, training for 30 rounds with 5 local epochs per round. Models use a three-layer fully-connected architecture with ReLU activations and an evidential output layer for Murmura. We use SGD with learning rate 0.01 and batch size 32. For evidential trust aggregation, we set self-weight $\alpha=0.5$, accuracy weight $w_{a}=0.5$, and initial trust threshold $\tau=0.3$. All experiments use 5 random seeds; we report mean accuracy and standard deviation across nodes.
@@ 2. RQ2:
    Does uncertainty-based peer selection improve convergence speed compared to baseline aggregation strategies?
@@ VII-D RQ2: Convergence Analysis
The fast convergence has practical implications for wearable IoT deployments: fewer communication rounds reduce energy consumption and network bandwidth usage, critical constraints for battery-powered devices. Combined with the single-pass uncertainty computation, Murmura provides an efficient path to personalized models.
@@ VII-E RQ3: Hyperparameter Sensitivity
Trust threshold ($\tau$) 98.4 $\pm$ 0.2 97.3 $\pm$ 0.2 78.9 $\pm$ 0.2
Vacuity threshold ($\nu$) 98.4 $\pm$ 0.2 97.2 $\pm$ 0.1 78.7 $\pm$ 0.1
The results demonstrate remarkable stability across hyperparameter choices. Standard deviations remain below 1% for UCI HAR and PAMAP2 across all parameters, and below 4% for the more challenging PPG-DaLiA dataset. The accuracy weight ($\lambda\in\{0.3,0.5,0.7,0.9\}$) balances prediction accuracy against uncertainty in trust computation; performance remains stable across this range. The self-weight ($\omega\in\{0.3,0.5,0.6,0.7,0.9\}$) controls personalization strength; higher values favor local models while lower values enable more collaboration. The trust threshold ($\tau\in\{0.05,0.1,0.2,0.3\}$) and vacuity threshold ($\nu\in\{0.3,0.5,0.7,0.9\}$) govern peer filtering stringency; the framework remains robust across a wide range. This stability reduces the need for extensive hyperparameter tuning in deployment, an important practical consideration for wearable IoT applications where per-device optimization is infeasible.
@@ VII-F Discussion
The results on PPG-DaLiA, while showing improvement over baselines at moderate heterogeneity, reveal limitations under extreme heterogeneity. This dataset exhibits substantial inter-subject variability in physiological signals, creating scenarios where few truly compatible peers exist. In such cases, Murmura’s conservative filtering leads to near-isolated training, which, while preserving personalization, limits knowledge transfer. Future work could explore graduated trust mechanisms that extract partial value from moderately compatible peers.
VIII Conclusion and Future Work
We presented Murmura, a framework enabling trust-aware model personalization in decentralized federated learning through evidential uncertainty quantification. The key insight is that epistemic uncertainty from evidential models directly indicates peer compatibility: high uncertainty when evaluating peer models on local data reveals distributional mismatch, enabling nodes to maintain personalized models through selective collaboration rather than forced consensus.
[…]
Future work should investigate temporal dynamics where peer compatibility may evolve over time, enabling adaptive personalization in changing environments. Deployment on physical wearable devices would validate practical applicability and reveal real-world performance characteristics. Additionally, graduated trust mechanisms that extract partial value from moderately compatible peers could improve performance when few highly compatible peers exist.
Murmura demonstrates that evidential deep learning provides a principled foundation for compatibility-aware personalization in decentralized settings, enabling collaborative learning that respects the natural heterogeneity of distributed data while achieving significant improvements in convergence speed and personalization quality.
Code Availability
The complete implementation of Murmura is available as open-source software at https://github.com/Cloudslab/murmura.

### v2

@@ VII-A1 Datasets
  • UCI HAR [1]: Smartphone accelerometer and gyroscope data from 30 subjects performing 6 activities (walking, walking upstairs, walking downstairs, sitting, standing, lying). Contains 10,299 samples with 561 extracted features.
  • PAMAP2 [17]: IMU data from 9 subjects performing 12 activities including household and exercise tasks. We extract 40 features per sliding window from three body-worn sensors.
  • PPG-DaLiA [16]: Photoplethysmography and accelerometer signals from 15 subjects under real-life conditions. We formulate activity classification from 8 activity types with time-frequency features.
@@ VII-A2 Baselines
  • FedAvg [13]: Standard federated averaging without Byzantine resilience or personalization mechanisms.
  • BALANCE [6]: Distance-based filtering with adaptive thresholds for Byzantine resilience and personalization.
  • Sketchguard [15]: Count-Sketch compression with L2 distance based filtering for communication-efficient DFL.
  • UBAR [7]: Two-stage filtering combining distance-based screening with loss-based evaluation.
@@ VII-A3 Implementation Details
We deploy 30 nodes in a fully-connected topology, training for 30 rounds with 5 local epochs per round. Models use a three-layer fully-connected architecture with ReLU activations and an evidential output layer for Murmura. We use SGD with learning rate 0.01 and batch size 32. For evidential trust aggregation hyperparameters, we set: self-weight $\omega=0.5$ to balance personalization and collaboration, accuracy weight $w_{a}=0.5$ to equally weight prediction correctness and uncertainty, initial trust threshold $\tau_{\min}^{(0)}=0.3$, adaptive tightening parameters $\gamma_{\tau}=0.5$ and $\kappa=1.0$, and epistemic uncertainty threshold $\tau_{u}=0.7$ to filter high-uncertainty peers. All experiments use 5 random seeds; we report mean results across nodes.
@@ 2. RQ2:
    Does uncertainty-based peer selection improve convergence speed compared to baseline strategies?
@@ VII-E RQ3: Hyperparameter Sensitivity
Trust threshold ($\tau_{\min}^{(0)}$) 98.4 $\pm$ 0.2 97.3 $\pm$ 0.2 78.9 $\pm$ 0.2
Uncertainty threshold ($\tau_{u}$) 98.4 $\pm$ 0.2 97.2 $\pm$ 0.1 78.7 $\pm$ 0.1
The results demonstrate remarkable stability across hyperparameter choices. Standard deviations remain below 1% for UCI HAR and PAMAP2 across all parameters, and below 4% for the more challenging PPG-DaLiA dataset. The accuracy weight ($w_{a}\in\{0.3,0.5,0.7,0.9\}$) balances prediction accuracy against uncertainty in trust computation; performance remains stable across this range. The self-weight ($\omega\in\{0.3,0.5,0.6,0.7,0.9\}$) controls personalization strength; higher values favor local models while lower values enable more collaboration. The trust threshold ($\tau_{\min}^{(0)}\in\{0.05,0.1,0.2,0.3\}$) and uncertainty threshold ($\tau_{u}\in\{0.3,0.5,0.7,0.9\}$) govern peer filtering stringency; the framework remains robust across a wide range. This stability reduces the need for extensive hyperparameter tuning in deployment, an important practical consideration for wearable IoT applications where per-device optimization is infeasible.
@@ VII-F Discussion
The results on PPG-DaLiA, while showing improvement over baselines at moderate heterogeneity, reveal limitations under extreme heterogeneity. This dataset exhibits substantial inter-subject variability in physiological signals, creating scenarios where few truly compatible peers exist. In such cases, Murmura’s conservative filtering leads to near-isolated training, which, while preserving personalization, limits knowledge transfer. Future work could explore graduated trust mechanisms that extract value from moderately compatible peers.
VIII Conclusions and Future Work
We presented Murmura, a framework enabling trust-aware model personalization in DFL through evidential uncertainty quantification. The key insight is that epistemic uncertainty directly indicates peer compatibility: high uncertainty when evaluating peer models on local data reveals distributional mismatch, enabling selective collaboration rather than forced consensus. Experimental evaluation on three wearable IoT datasets demonstrates that Murmura reduces performance degradation from IID to non-IID conditions by a factor of 20 compared to FedAvg (0.9% vs 19.3%), converges 7.4 times faster, and maintains stable accuracy across hyperparameter choices. Future work should investigate temporal dynamics where peer compatibility evolves over time, deployment on physical wearable devices to validate practical applicability, and graduated trust mechanisms that extract partial value from moderately compatible peers when few highly compatible peers exist. These results establish EDL as a principled foundation for compatibility-aware personalization in decentralized heterogeneous environments.